The force required to flex a nasal strip and apply it to the nose is 0.23 N; the energy stored in the strip when flexed is 0.0025 J. Assume the strip to be an ideal spring for the following questions. A.) Find the distance through which the strip is flexed.
Answer #1

F = 0.23N , U = 0.0025J

Force due to spring is F = kx

And energy stored in spring is U = 1/2 kx^2

=) U = 1/2 Fx

=) x = 2U/F = 2×0.0025/0.23 = 0.02174m = 2.17cm

Required distance x = 2.17cm

Note : in case it will show incorrect please try simple form of distance = U/F = .0.0025/0.23 = 0.01087m = 1.08cm.